Country Financial, a financial services company, uses surveys of adults age 18 and older to determine if personal financial fANSWER
Answer #1

b) The null and alternative hypothesis are: Ho: P.-P₂ = 0 Haip, -P₂70 This is a Two Failed Test. The sample proportion indicaThe rest statistic in, Z = PCI-E) ( + + - + 0.410 – 0.35 26.3816) (1-0.3816) (1000 + 2 ) = 2.6881703 = 2.69 P-value = 2* Plz>

> The only mistake there is that, the second endpoint should be .1037 instead of .1307.

tom1113 Thu, Nov 4, 2021 4:40 PM
Answer #2

The P-Value is .0072

source: Taxprep excell sheet provided by cengage
answered by: Caleb
Answer #3

my answer was correct at p value 0.0071 though however, upper interval was 0.1037 so (.0163 to 0.1037)

answered by: Yr